# Transit Network Design with Two-Level Demand Uncertainties: A Machine Learning and Contextual Stochastic Optimization Framework

### v1

@@ 4.3 Experiments: Choice Modeling and Evaluating 2LRC-TND
The second group of experiments evaluates the 2LRC-TND framework by comparing it against two benchmarks. The first benchmark, Fixed-Demand (FD), designs a network using only the 6,659 existing transit riders in a deterministic manner, without considering demand uncertainties. The second benchmark, Naive Rule-Based (RB), with no involvement of Machine Learning, uses hand-crafted rules to perform the same functions as $\mathcal{C}_{\mathrm{core}}$ and $\mathcal{C}_{\mathrm{adopt}}$. A simple rule-based model example is: "If both the origin and destination of a trip are within a 5-minute walk distance to a transit station, assign a core trip probability of 0.8; otherwise, assign 0.2." In general, a rule-based approach requires extensive domain expertise, but the design of such rules is beyond the scope of this paper. For each of these runs (2LRC-TND and Benchmarks), two budget levels are tested. These values are chosen based on the total cost of activating all available arcs from 8AM to 9AM, which is approximately \$58K. A budget of \$30K represents roughly half of this total, while \$35K corresponds to 60%, allowing for a meaningful comparison under different budget constraints.
@@ 5.2 Results on 2LRC-TND
Table 3 summarizes the computational results on optimization. Random Forest is employed as $\mathcal{C}_{\mathrm{core}}$ and $\mathcal{C}_{adopt}$ in 2LRC-TND because its simplicity for implementation (compared to DNNs) and superior performance in prediction results demonstrated in prior experiments. Selected visualizations (due to page limit) of designs are shown in Figures 4. Examining the network designs, it appears that all runs open a similar number of arcs within the same budget. This outcome occurs because the model tries to utilize the entire available budget.
The second benchmark, Naive Rule-based (RB), uses nested if-else conditions based on travel time and origin-destination locations to generate probabilities, serving the same purpose as $\mathcal{C}_{\mathrm{core}}$ and $\mathcal{C}_{\mathrm{adopt}}$ in 2LRC-TND but without machine learning. The first observation is, using the rule-based model, the CP can be solved optimally within a short time, whereas 2LFC-TND ($B$ = 30K) only reports the best feasible solution after the run time limit of 12 hours is reached. This is likely due to the machine learning model predicting a greater number of core riders, which requires the CP model to provide a path for each core rider, thereby increasing computational difficulties. Secondly, although the core riders differ across runs, all experiments demonstrate that 2LRC-TND achieves a notable number of adopted riders. Specifically, RF yields a higher adoption rate by learning directly from data, while Naive RB depends on manually defined rules that lack consideration of geographical and demographic information.
@@
Naive RB 1.98 3319 2139 36040 6182 (17.15%) 8321 8819 57.03 8313.00 (0)
2LFC-TND 3.88 3313 4259 33920 11745 (34.63%) 16005 16655 56.56 15999.53 (0)
@@
Lastly, the designs are evaluated using 1000 ($I^{\prime}$) additional SAA scenarios. Results show that all core riders are served under the respective designs with no violation, which is likely because each $\mathcal{P}_{r}$ contains four paths and the fixed rail arcs are effectively utilized by core riders. Furthermore, the $\mathbf{Eval(\mathbf{z})}$ results closely match the coverage values from the evaluation with $I$, indicating that the results are highly stable after solving from only 50 SAA scenarios.
@@ 6 Conclusion
This paper introduces the 2LRC-TND framework to address the limitations of traditional transit network design approaches that assume fixed demand. Leveraging CSO, the framework integrates machine learning-based choice models and constraint programming-based optimization models to produce transit network designs. A large-scale case study over 38,000 trips from Atlanta demonstrates that 2LRC-TND can improve ridership within reasonable budget limits, highlighting its practicality and potential social impact. Several promising directions remain for future research including design of more targeted surveys for enhancing data collection and extension of 2LRC-TND to emerging transportation systems.

### v2

@@ 4.3 Experiments: Choice Modeling and Evaluating 2LRC-TND
The second group of experiments evaluates the 2LRC-TND framework by comparing it against three benchmarks. The first benchmark, Fixed-Demand (FD), designs a network using only the 6,659 existing transit riders in a deterministic manner, without considering demand uncertainties. The second benchmark, Naive Rule-Based (RB), with no involvement of Machine Learning, uses hand-crafted rules to perform the same functions as $\mathcal{C}_{\mathrm{core}}$ and $\mathcal{C}_{\mathrm{adopt}}$. A simple rule-based model example is: "If both the origin and destination of a trip are within a 5-minute walk distance to a transit station, assign a core trip probability of 0.8; otherwise, assign 0.2." In general, a rule-based approach requires extensive domain expertise, but the design of such rules is beyond the scope of this paper. The third benchmark, Deterministic, employs the same machine learning models as 2LRC-TND but considers them under a fixed threshold of 0.5 to obtain deterministic decisions and designs the network under this setting.
For each of these runs (2LRC-TND and Benchmarks), two budget levels are tested. These values are chosen based on the total cost of activating all available arcs from 8AM to 9AM, which is approximately \$58K. A budget of \$30K represents roughly half of this total, while \$35K corresponds to 60%, allowing for a meaningful comparison under different budget constraints.
@@ 5.2 Results on 2LRC-TND
Table 3 summarizes the computational results on optimization. Random Forest is employed as $\mathcal{C}_{\mathrm{core}}$ and $\mathcal{C}_{adopt}$ in 2LRC-TND because its simplicity for implementation (compared to DNNs) and superior performance in prediction results demonstrated in prior experiments. Selected visualizations (due to page limit) of designs are shown in Figures 4. Examining the network designs, it appears that all runs open a similar number of arcs within the same budget. This outcome occurs because the model tries to utilize the entire available budget. The resulting networks largely preserve existing MARTA corridors while reallocating capacity toward underserved areas with high latent demand.
@@
Naive RB 1.98 3319 2139 36040 6182 (17.15%) 8321 8819 57.03 8313.00 0(0)
Deterministic 0.03 2838 4776 33403 966 0(2.89%) 5742 5742 53.12 13474.61 (15)
2LRC-TND 3.88 3313 4259 33920 11745 (34.63%) 16005 16655 56.56 15999.53 0(0)
The second benchmark, Naive Rule-based (RB), uses nested if-else conditions based on travel time and origin-destination locations to generate probabilities, serving the same purpose as $\mathcal{C}_{\mathrm{core}}$ and $\mathcal{C}_{\mathrm{adopt}}$ in 2LRC-TND but without machine learning. The first observation is, using the rule-based models, the CP can be solved optimally within a short time, whereas 2LRC-TND ($B$ = 30K) only reports the best feasible solution after the run time limit of 12 hours is reached. This is likely due to the machine learning model predicting a greater number of core riders, which requires the CP model to provide a path for each core rider, thereby increasing computational difficulties. Secondly, although the core riders differ across runs, all experiments demonstrate that 2LRC-TND achieves a notable number of adopted riders. Specifically, RF yields a higher adoption rate by learning directly from data, while Naive RB depends on manually defined rules that lack consideration of geographical and demographic information.
To demonstrate the importance of stochastic modeling, a third benchmark employing deterministic choice models is evaluated against 2LRC-TND. This configuration utilizes the same machine learning models as 2LRC-TND; however, they generate deterministic binary decisions using a typical fixed threshold of 0.5, thus eliminating the need for SAA scenarios in $I$. In this case study, since the predicted probability of adopting a path for latent trips is typically below 0.5 in most cases, this deterministic approach underestimates potential ridership adoption. As shown in Table 3, the deterministic approach achieves a significantly weaker adoption rate across both budget levels.
@@
Lastly, the designs are evaluated using 1000 ($I^{\prime}$) additional SAA scenarios. Results show that all core riders are served under the respective designs with no violation in most of the cases, which is likely because each $\mathcal{P}_{r}$ contains four paths and the fixed rail arcs are effectively utilized by core riders. However, Deterministic case exhibits 15 constraint violations when evaluated on these scenarios indicating the need for the stochastic decision-making models. Additionally, the $\mathbf{Eval(\mathbf{z})}$ results closely match the coverage values from the evaluation with $I$ for the 2LRC-TND approach, highlighting that the results are highly stable after solving from only 50 SAA scenarios.
@@ 6 Conclusion
This paper introduces the 2LRC-TND framework to address the limitations of the traditional transit network design approaches that assume fixed demand. Leveraging CSO, the framework integrates machine learning-based choice models and constraint programming-based optimization models to produce transit network designs. A large-scale case study over 38,000 trips from Atlanta demonstrates that 2LRC-TND can improve ridership within reasonable budget limits, highlighting its practicality and potential social impact. Several promising directions for future research include design of more targeted surveys for enhancing data collection and extension of 2LRC-TND to emerging transportation systems.